# DyRoNet: Dynamic Routing and Low-Rank Adapters for Autonomous Driving Streaming Perception

### v3

@@ 5 Conclusion
In conclusion, we present the Dynamic Routering Network (DyRoNet), a system that dynamically selects specialized detectors for varied environmental conditions with minimal computational overhead. Our innovative increase-boosting fine-tuning, featuring a Low-Rank Adapter, mitigates distribution bias and overfitting, enhancing scene-specific performance. Experimental results validate DyRoNet’s state-of-the-art performance, offering a benchmark for streaming perception and insights for future research. In the future, DyRoNet’s principles will undoubtedly inform the development of more advanced, reliable systems.
@@ Appendix A Factor Analysis in Streaming Perception
In development of DyRoNet, we undertook an extensive survey and analysis to identify key influencing factors in autonomous driving scenarios that could potentially impact streaming perception. This analysis utilized the Argoverse-HD dataset Li et al. [2020], a benchmark in the field of streaming perception. The primary goal of this factor analysis was to isolate the most critical factor affecting streaming perception performance. As elaborated in the main text, our comprehensive analysis led to the identification of the speed of the environment as the predominant factor. Consequently, DyRoNet is tailored to address this specific aspect. Our analysis focuses on four primary elements: weather conditions, object quantity, small object proportion, and environmental speed. We methodically examined each of these factors to evaluate their respective impacts on streaming perception within autonomous driving.
@@ A.1 Impact of Weather Conditions
To evaluate the impact of weather conditions on streaming perception, we conducted tests using a range of pre-trained models from StreamYOLO Yang et al. [2022], LongShortNet Li et al. [2023], and DAMO-StreamNet He et al. [2023], employing various scales and settings. The results, presented in Tab. 8, indicate that performance is generally better during Day conditions compared to Night. This confirms that weather conditions indeed influence streaming perception.
@@
Table 8: Offline Evaluation Results on the Argoverse-HD Validation Dataset: It records the sAP scores across the 0.50 to 0.95 range for each clip. The optimal and worst results are highlighted in green and red font under the same weather conditions. The notation “l high” is used as an abbreviation for the resolution 1200x1920, providing a concise representation of the data.
@@ A.3 Analysis of the Proportion of Small Objects
The influence of small objects on perception models, particularly in autonomous driving scenarios, has been underscored in studies like Mahaur and Mishra [2023] and Yang et al. [2022]. In such scenarios, even minor shifts in viewing angles can cause notable relative displacement of small objects, posing a challenge for perception models in processing streaming data effectively. This observation prompted us to closely examine the proportion of small objects in the environment.
@@ C.1 Pre-trained Model Selection
As outlined in the main paper, our implementation of DyRoNet incorporates three existing models as branches within the Model Bank $\mathcal{P}$: StreamYOLOYang et al. [2022], LongShortNetLi et al. [2023], and DAMO-StreamNetHe et al. [2023]. These models were selected due to their specialized features and proven effectiveness in streaming perception tasks. StreamYOLO is unique for its two additional pre-trained weight variants, each tailored for different streaming processing speeds. This feature allows for adaptable performance depending on the speed requirements of the streaming task. In contrast, LongShortNet and DAMO-StreamNet are equipped with pre-trained weights optimized for high-resolution image processing, making them suitable for scenarios where image clarity is paramount.
To ensure a diverse and versatile range of options within the Model Bank, our implementation of DyRoNet selectively utilizes the Small (S), Medium (M), and Large (L) variants of the pre-trained weights from each model. This choice enables a balanced mix of processing speeds and resolution handling capabilities, catering to a wide range of streaming perception scenarios. The specific details regarding the number of parameters for these pre-trained models can be found in Table 12, which provides a comparative overview to help in understanding the computational complexity for different tasks.
@@ C.2 Setting of Hyperparameters
Acknowledgments
Zhi-Qi Cheng’s research in this project was supported by the US Department of Transportation, Office of the Assistant Secretary for Research and Technology, under the University Transportation Center Program (Federal Grant Number 69A3551747111), and Intel and IBM Fellowships.
Contribution Statement
Xiang Huang, Jun-Yan He, and Zhi-Qi Cheng worked together on the research idea and were fully involved in the experimental work. Jun-Yan He and Zhi-Qi Cheng made revisions to the manuscript, offering valuable insights and recommendations for the experimental procedures. The manuscript was then reviewed by Chengyang Li, Wangmeng Xiang, Baigui Sun, and Xiao Wu. Zhi-Qi Cheng served as the corresponding author and led the entire project.

### v4

@@ 5 Conclusion
In conclusion, we present the Dynamic Routering Network (DyRoNet), a system that dynamically selects specialized detectors for varied environmental conditions with minimal computational overhead. Our Low-Rank Adapter mitigates distribution bias, thereby enhancing scene-specific performance. Experimental results validate DyRoNet’s state-of-the-art performance, offering a benchmark for streaming perception and insights for future research. In the future, DyRoNet’s principles will inform the development of more advanced, reliable systems.
Acknowledgments
Zhi-Qi Cheng’s research in this project was supported in part by the US Department of Transportation, Office of the Assistant Secretary for Research and Technology, under the University Transportation Center Program (Federal Grant Number 69A3551747111), and Intel and IBM Fellowships.
@@ Appendix A Factor Analysis in Streaming Perception
In development of DyRoNet, we undertook an extensive survey and analysis to identify key influencing factors in autonomous driving scenarios that could potentially impact streaming perception. This analysis utilized the Argoverse-HD dataset [20], a benchmark in the field of streaming perception. The primary goal of this factor analysis was to isolate the most critical factor affecting streaming perception performance. As elaborated in the main text, our comprehensive analysis led to the identification of the speed of the environment as the predominant factor. Consequently, DyRoNet is tailored to address this specific aspect. Our analysis focuses on four primary elements: weather conditions, object quantity, small object proportion, and environmental speed. We methodically examined each of these factors to evaluate their respective impacts on streaming perception within autonomous driving.
@@ A.1 Impact of Weather Conditions
To evaluate the impact of weather conditions on streaming perception, we conducted tests using a range of pre-trained models from StreamYOLO [35], LongShortNet [19], and DAMO-StreamNet [10], employing various scales and settings. The results, presented in Tab. 10, indicate that performance is generally better during Day conditions compared to Night. This confirms that weather conditions indeed influence streaming perception.
@@
Table 10: Offline Evaluation Results on the Argoverse-HD Validation Dataset: It records the sAP scores across the 0.50 to 0.95 range for each clip. The optimal and worst results are highlighted in green and red font under the same weather conditions. The notation “l high” is used as an abbreviation for the resolution $1200\times 1920$, providing a concise representation of the data.
@@ A.3 Analysis of the Proportion of Small Objects
The influence of small objects on perception models, particularly in autonomous driving scenarios, has been underscored in studies like [22] and [35]. In such scenarios, even minor shifts in viewing angles can cause notable relative displacement of small objects, posing a challenge for perception models in processing streaming data effectively. This observation prompted us to closely examine the proportion of small objects in the environment.
@@ B.1 Inference Time Analysis
B.2 Statistic of model selection
We also provide statistics on DyRoNet’s selection of different models during both training and inference time in Tab.14. From the results, it can be observed that DAMO-StreamNet (M+L) exhibits a bias to select the second model during inference time, leading to a similar performance as DAMO-StreamNet L. However, under normal circumstances, DyRoNet can still dynamically choose the appropriate model based on input conditions.
Model training time inference time
Combination Model 1 Model 2 Model 1 Model 2
SYOLO S+M 14.24% 85.76% 5.95% 94.05%
SYOLO S+L 10.98% 89.02% 4.83% 95.17%
SYOLO M+L 37.53% 62.47% 94.67% 5.33%
LSN S+M 13.05% 86.95% 81.65% 18.35%
LSN S+L 7.28% 92.72% 17.26% 82.74%
LSN M+L 30.86% 69.14% 19.87% 80.13%
DAMO S+M 6.26% 93.74% 0.00% 100.00%
DAMO S+L 35.29% 64.71% 3.69% 96.31%
DAMO M+L 84.61% 15.39% 0.02% 99.98%
Table 14: The statistics of model selection by DyRoNet under different model choices during both training and inference time.
B.3 The comparison between Speed Router and $\mathbb{E}[\Delta I_{t}]$
Model Bank $\mathbb{E}[\Delta I_{t}]$ (sAP) Speed Router (sAP)
$\text{StreamYOLO}_{\text{S + M}}$ 31.5 32.6 (+1.1)
$\text{StreamYOLO}_{\text{S + L}}$ 32.9 35.0 (+2.1)
$\text{StreamYOLO}_{\text{M + L}}$ 34.2 34.6 (+0.4)
Table 15: Comparison of Speed Router and $\mathbb{E}[\Delta I_{t}]$. Where $\mathbb{E}[\Delta I_{t}]$ means directly select model by the sign of $\mathbb{E}[\Delta I_{t}]$ without using Speed Router.
We also consider a special case id Tab.15, where the model selection only base on the mean of $\Delta I_{t}$ without using Speed Router, which is denoted as $\mathbb{E}[\Delta I_{t}]$. To be specific, the larger model is selected when $\mathbb{E}[\Delta I_{t}]>0$ and minor model is selected otherwise. Unlike Tab.5 in the main text, both methods here are trained for 5 epoch using LoRA fine-tuning. From the results in Tab.15, it can be seen that our proposed Speed Router has significant advantages compared to directly using $\mathbb{E}[\Delta I_{t}]$ to select branches.
$=0$ $>0$ $<0$
train 0.24% 48.22% 51.55%
test 0.30% 49.85% 49.85%
val 0.17% 49.18% 50.66%
Table 16: Statistics of the sign of $\mathbb{E}(\Delta I_{t})$ over Argoverse-HD.
Furthermore, in Tab.16, we also conducted the statistic the sign of $\mathbb{E}[\Delta I_{t}]$ on the Argoverse-HD. Results with absolute values less than 1e-6 were considered equal to 0. The results reveal that evenly distributing training across models did not effectively adapt them to varying speeds as our Speed Router did.
@@ C.1 Pre-trained Model Selection
As outlined in the main paper, our implementation of DyRoNet incorporates three existing models as branches within the Model Bank $\mathcal{P}$: StreamYOLO[35], LongShortNet[19], and DAMO-StreamNet[10]. These models were selected due to their specialized features and proven effectiveness in streaming perception tasks. StreamYOLO is unique for its two additional pre-trained weight variants, each tailored for different streaming processing speeds. This feature allows for adaptable performance depending on the speed requirements of the streaming task. In contrast, LongShortNet and DAMO-StreamNet are equipped with pre-trained weights optimized for high-resolution image processing, making them suitable for scenarios where image clarity is paramount.
To ensure a diverse and versatile range of options within the Model Bank, our implementation of DyRoNet selectively utilizes the Small (S), Medium (M), and Large (L) variants of the pre-trained weights from each model. This choice enables a balanced mix of processing speeds and resolution handling capabilities, catering to a wide range of streaming perception scenarios. The specific details regarding the number of parameters for these pre-trained models can be found in Tab.17, which provides a comparative overview to help in understanding the computational complexity for different tasks.
@@ C.2 Setting of Hyperparameters
Appendix D Details of experiment on NuScenes-H dataset
To meet the requirements of streaming perception tasks, nuScenes-H [30] enhances the commonly used autonomous driving perception dataset nuScenes [2] by increasing the annotation frequency from 2Hz to 12Hz. While nuScenes encompasses data from three modalities—Camera, LiDAR, and Radar—nuScenes-H provides dense 3D object annotations exclusively for the 6 sensors of Camera modality.
As mentioned in the main text, we trained and evaluated DyRoNet on the nuScenes-H dataset. To accommodate the requirements for 2D object detection, the 3D object annotations in nuScenes-H are converted to 2D using publicly available conversion scripts. All experiments were conducted exclusively using the CAM_FRONT viewpoint. The dataset partition details are summarized in Tab. 18, which includes the number of video clips, video frames, and the instance counts for each object category within the subsets. As it shows in Tab. 18, limited or even absent annotation for some categories resulted in lower overall test performance. For clarity, this dataset is referred to as nuScenes-H 2D as follows.
Before training DyRoNet, YOLOX [6] was trained for 80 epochs on nuScenes-H 2D to obtain pretrained weights. These weights were then used to initialize the branch models within DyRoNet. During the training of DyRoNet, each individual branch was trained for 10 epochs, followed by 5 epochs of training for the router. All other training settings were consistent with those described in the main text. As indicated by the experimental results presented, DyRoNet maintains strong selection capabilities across different branches on other datasets, demonstrating its adaptability under practical application conditions.
train set test set
# of video clips 120 30
# of frames 26705 6697
# of anno 225346 71819
adult 32200 13920
child 22 142
wheelchair 0 0
stroller 0 174
personal_mobility 0 2
police_officer 0 0
construction_worker 1573 362
animal 22 0
car 100487 25356
motorcycle 4958 330
bicycle 1844 1248
bus.bendy 531 283
bus.rigid 4854 1161
truck 21801 4934
construction 2154 1200
emergency.ambulance 61 0
emergency.police 112 0
trailer 6799 805
barrier 33058 10568
trafficcone 8654 10096
pushable_pullable 5191 649
debris 666 348
bicycle_rack 359 241
Table 18: Dataset partition of nuScenes-H 2D, includes the number of video clips (# of video clip), video frames (# of video clip), and the instance counts for each object category (# of anno) within the subsets.